OpenStudy (anonymous):

sketch a graph for the function; f(x)=1/2(x-|x|)

OpenStudy (tkhunny):

For x > 0, |x| = x and x - |x| = x - x = 0 What about x < 0?

OpenStudy (anonymous):

hmm....I am not sure

OpenStudy (tkhunny):

You'll have to do better than that. For x < 0, |x| = -x Do you believe this?

OpenStudy (anonymous):

lt's the opposite when it is less

OpenStudy (tkhunny):

|-3| = -(-3) = +3 |-6| = -(-6) = +6 You must be certain that this is so. For x < 0, |x| = -x.

OpenStudy (anonymous):

(x-|x|) the x-value must be less than zero. That means when you're sketching the curve, you now know that the positive side of the x-axis cannot be used in this instance.

OpenStudy (tkhunny):

?? The Domain is \(\mathbb{R}\). There is no restriction. For \(x \ge 0\), |x| = x and we have x - |x| = x - (x) = 0 For \(x < 0\), |x| = -x and we have x - |x| = x - (-x) = x + x = 2x

OpenStudy (anonymous):

No, I don't think so.I'm sure of it that if x<0, you would have a negative denominator.

OpenStudy (anonymous):

For x<0 , (-x) - |(-x)|=-x-(x)=-x-x=-2x

OpenStudy (tkhunny):

The only denominator is 2. @malibugranprix2000, do you mean \(\dfrac{1}{2}(x - |x|)\) as you have written, or did you intend \(\dfrac{1}{2(x-|x|)}\), which is not what you have written?

OpenStudy (tkhunny):

No, x is still x, no matter the sign of x.

OpenStudy (anonymous):

No there's no absolute value sign on the x that's free.

OpenStudy (anonymous):

If x was negative. you would have -2x regardless of anything weird happening.

OpenStudy (tkhunny):

No, if x is negative, then 2x would be a negative number. Why would it suddenly change to positive (-2x)?

OpenStudy (anonymous):

What are you on about? What you said was practically the same. negative 2x is the same as positive (-2x)....-2x=+(-2x)?

OpenStudy (anonymous):

@malibugranprix2000 Please draw your correct equation.

OpenStudy (tkhunny):

No, that is just wrong. For x < 0, 2x < 0 and -2x > 0

OpenStudy (anonymous):

Ah okay. Yep, you're right. Sorry about that mate.

OpenStudy (anonymous):

It is 1/2(x-|x|) not 1/2....

OpenStudy (anonymous):

it is not 1 over 2(x-|x|)

OpenStudy (anonymous):

Okay. I think tkhunny's got it here then.

OpenStudy (tkhunny):

You wrote the same thing, which is \(\dfrac{1}{2}(x - |x|)\). Please refer to this as #1. If you mean \(\dfrac{1}{2(x-|x|)}\), please consider the Order of Operations and use more parentheses. Please refer to this as #2. Which is it? Many thanks to @Azteck for pointing out the possible ambiguity.

OpenStudy (anonymous):

number 1

OpenStudy (tkhunny):

For #1, we have: For x≥0 , |x| = x and we have x - |x| = x - (x) = 0 an horizontal line on the non-negative x-axis. For x<0 , |x| = -x and we have x - |x| = x - (-x) = x + x = 2x a line descending from the Origin. For #2, we have For x≥0 , |x| = x and we have x - |x| = x - (x) = 0 and this is not in the Domain. For x<0 , |x| = -x and we have x - |x| = x - (-x) = x + x = 2x and we have half an hyperbola. An interesting problem, either way.

OpenStudy (anonymous):

thanks @tkhunny and @Azteck for the explanation